OpenStudy (anonymous):

The sum of two numbers, x and y, is 12. The difference of x and two times y is 6. What are the values of x and y?

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

Write the equations: The sum of two numbers, x and y, is 12: \[x+y = 12\]The difference of x and two times y is 6:\[x-2y=6\] Now you have a system of two equations in two unknowns. Do you know how to solve it?

OpenStudy (anonymous):

I do not sir

OpenStudy (whpalmer4):

Okay, there are a number of ways to do so, but two of the easiest are substitution and elimination. Let's do substitution. Can you solve the first equation \((x+y=12\)) for either \(x\) or \(y\)? Your choice.

OpenStudy (whpalmer4):

Okay, we'll solve for \(x\). \[x+y=12\]Subtract \(y\) from both sides\[x+y-y=12-y\]\[x=12-y\]Was that so difficult? Now, in the second equation, \(x-2y=6\), you need to replace \(x\) with \((12-y)\)wherever you see it. That will give you an equation in terms of \(y\) only. Solve that, and you'll have your value of \(y\). Then use it in \(x = 12-y\) to find \(x\).

OpenStudy (anonymous):

x=12-y ?

OpenStudy (anonymous):

10?

OpenStudy (whpalmer4):

10? What is 10? The number of pizzas you're going to order for me?

OpenStudy (jdoe0001):

better be small-size pizzas, I could only eat one family-size myself =), couldn't eat 2 though

OpenStudy (whpalmer4):

So, what are the answers to this problem?

OpenStudy (whpalmer4):

@jdoe0001 I'll shoot you a PM when the pizzas arrive, you can come help me eat them :-)